OpenStudy (anonymous):

Let n(A u B)=50 , n(B)=20, n(A)=45, and U=A u B.

OpenStudy (anonymous):

you meant P(A∩B)? n(A U B) = n(A) + n(B) - n(A∩B) 50 = 45 + 20 - n(A∩B) you can do the math

OpenStudy (raden):

what's the question here ?

OpenStudy (anonymous):

Find (A` u B`)

OpenStudy (anonymous):

(A' U B') = (A∩B)' = U - (A∩B) 50 - 15 = 35